Solve the system of equations and express the answer as an ordered
pair of the form (x,y) or write NO SOLUTION if applicable:
x + 2y = 5
2x - 2y = 4

Answer by ewatrrr(24785)

 
Hi,
Solving the system of equations by the elimination method
x + 2y = 5 Green line
2x - 2y = 4 Blue line
3x = 9
x = 3 an y = 1 ((5-3)/2
Ordered pair (3,1) the solution for this system